Tech Staff Meeting Notes January 5, 2016

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights


Put AO/TA conversions on hold for now. Jim requested the extra sim card now so that it could be used for several noise studies on deck. The six stations in the AO network are being decommissioned and the hardware returned to the ANSS Depot. CERI offered to take over O&M for two: U40A (near Mountain Home, AR) and X40A (east of Hot Springs). Jim did a recon on U40A and its doable but will take a little work. This effort will likely take about 6 months.

2015 Preventive Maintenance84% complete
NM BB/SM & SP: 67 done & 10 to go
NM SM: 9 done & 8 to go
AG BB/SM: 6 done & 0 to go
GPS: 13 done & 0 to go
Misc: 6 done & 2 to go

freewave serial reset installed at 10 of 11 broadbands. GLAT remains.


Taken from:

MORT - possible flooding

Misc Problems
NFAR under powered
CHNM -spikes see image
CPAR - spikes
HOVM - see image
NNAR - see image

Need to do at Recon at MORT to verify flooding; reconfigure to allow repeat of outer station TNMT until its dry enough to rebuild.

RVAR has slumped down the hill and the pipe vault is now bent. Need to install a repeater to get the outer station, MSAR, back online until access improves enough to rebuild RVAR.

All ISIS3 hardware except that slated for PLAR is installed. Have about 26 remaining, mostly MKTA and some NMAD outer stations.

RDST has a very large pole right next to the vault that will corrupt any data during strong ground motion. It needs to get moved; Dave has a couple options on the Rhodes campus (the name will change when its moved). We have a Basalt almost ready to go so we and install new rather than move old.

NODE Tasks

The win7 upgrades at nodes (mudworm is done). There are several win7 issues that Steve and James are working out before installing at remote locations.


The Wilder-SFTN DBB link conversion from aironets to the 2.5 GHz Ubiquiti Bullet-M radios was unsuccessful. They seem to work ok when connecting a pair of Ubiquities with a pair of Aironets but have networking issues when linked in with the two paths at SFTN. Next step is to try installing a switch at SFTN to help correctly route traffic

There are some inconsistencies in the configuratons of the aironets. Steve will review and sync configs.




This is much improved and may be removed from the list. The waveserver continues to have occasional very gappy periods. James found that rtpm was configured to store sms messages and there were an extremely large number of those files. They were cleaned up and reconfigured on all nodes. So far nmad seems to have improved and no significant gaps were produced over the weekend.

Gimli and bofur waveservers were reconfigured to run only a single set of waveservers on each. This has significantly improved performance.

AZTN/HDBR: Shahram submitted a proposal to TDOT to replace all the sensors with Episensors (including new cabinets, etc). We asked USGS to provide new radios so we can have a continuous buffer of all channels at AZTN. We'll then allow access to NSMP so they can extract event data. Unofficial word is TDOT approved the funding; waiting on signed contract.



Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Need to test farther away.


Data taken from:

Misc Problems
CPRT.HNN - see image
VHTN.EHE - see image
GMG - power problems
BCRT-EHN low amplitude

John's updated todo list .


Improve RCGA AC power and perform maintenance work on shed.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are t oo tall to get a signal out. Looking for an alternate site. Can we use a cell modem?


HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established.

Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.

GFM - Need to do a noise study for a closer site, then move it in the spring along with upgrading to S-13.


Replace VHTN E-W and N-S Instruments.

WMTN pit needs to be moved due to erosion.

We're still repeating CPRT through GTTN. The new internet host at the Corryton Senior Center for CPRT has been formally approved by Greg Hill the center coordinator and the Knox County facility supervisor.


No problems at this time.

AG Network

FCAR AC power upgrade was abandoned. A tree was removed improving sunlight Temporay additional solar was installed and seems to be working. Need to make it permanent.


Needs some clean up.

The FDL wiki is here:


OXF and LRAL are still down. Chris replaced the inverter and power brick for the modem but the station is still down. ASL crew is scheduled visit both sites in late January.

The seisnet group on umdrive is at You will need your uuid and password.

The seisnet wiki is at You will need your uuid and password.